Every week, Emily and Jacod each write one story for an online magazine. Emily is paid 200 dollars plus 5% of the advertising revenue for her story. Jacob is paid 150 dollars plus 10% of the advertising revenue for his story. If they both generated the same amount of advertising revenue, and received equal paychecks for the week, how much in dollars was each paid?
Found 3 solutions by fcabanski, mananth, stanbon:
Answer by fcabanski(1391)

Each is paid P = D + pr where D is some dollar amount, p is the percentage they earn of ad revenue, and r is the ad revenue.


For Emily that's P = 200+.05r and for Jacob it's P = 150+.1r


Note that r is the same in both cases, and the two are equal to each other (as given in the problem.


200+.05r=150+.1r


Subtract .05 r from both sides, and subtract 150 from both sides.


50 = .05r


Divide both sides by .05.


r = 1000


Substitute that r back in each equation to determine what they earned, and to prove the amounts are the same.


Emile = 200 + .05(1000) = 200 + 50 = 250


Jacob = 150 + .1(1000) = 150 + 100 = 250

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com



Answer by mananth(16949)

Emily --------------200 dollars plus 5%
Jacob---------------150+10%
let ad revenue be x
Emily = 200+0.05x
Jacob = 150 + 0.1x
Both are equal
200+0.05x=150+0.1x
200-150=0.1x-0.05x
0.05x=50
x=50/0.05
x=1000
Emily = 200+0.05x============> 200+0.05*1000=$250
Jacob = 150 + 0.1x===========>150+0.1*1000=$250
Both got $250
m.ananth@hotmail.ca

Answer by stanbon(75887)

Every week, Emily and Jacod each write one story for an online magazine. Emily is paid 200 dollars plus 5% of the advertising revenue for her story. Jacob is paid 150 dollars plus 10% of the advertising revenue for his story. If they both generated the same amount of advertising revenue, and received equal paychecks for the week, how much in dollars was each paid?
-----
Equation:
200 + 0.05a = 150 + 0.10a
0.05a = 50
5a = 5000
a = $1,000.00 (advertising revenue)
======================
Cheers,
Stan H.
==================